OpenStudy (anonymous):

Using disks or washers, find the volume of the solid obtained by rotating the region by the curves y=x^(2) and y=4 about the line y=4

OpenStudy (anonymous):

Using the disk method, do you see how the radius of each disk will be \[\text{radius}=4-x^2?\] The integral will then be \[\pi\int_{-2}^2(4-x^2)^2dx\] Taking advantage of symmetry, you have \[2\pi\int_{0}^2(4-x^2)^2dx\]

OpenStudy (anonymous):



OpenStudy (anonymous):